OpenStudy (anonymous):

Use the factor theorem to decide whether or not the second polynomial is a factor of the first: 6x^4+23x^3-4x^2+x+4; x+4

OpenStudy (anonymous):

\[6 x^4+23 x^3-4 x^2+x+4=(x+4) (2 x+1) \left(3 x^2-2 x+1\right) \]